表 1: W30cm×H30cm
光学系集光率反射回数
主鏡ファネル平均値最高値最低値壁ファネル
平面鏡  winston cone 24.0% 35%  15% 0.42  1.2
平面鏡 (tall)  winston cone 23.3% 32%  16% 0.44  1.2
平面鏡  box type  17.3% 23%  13% 0.48  1.6
振り上げ振り下ろし  winston cone 30.0% 35%  23% 0.40  1.1
振り上げ振り下ろし (tall) winston cone 32.0% 35% 28% 0.42  1.1
振り上げ振り下ろし  box type  20.3% 26%  16% 0.43  1.6
振り上げ振り下ろし  small cone  27.3% 36%  18% 0.38  1.0
振り上げ振り下ろし  large cone  29.0% 34%  24% 0.41  1.2
1 軸パラボラ  winston cone 25.6% 36%  11% 0.43  1.1
1 軸パラボラ (tall)  winston cone 25.2% 34%  12% 0.45  1.1
上げ下ろし (tall,n=1.05) winston cone 27.9% 35%  23% 0.44  1.1
上げ下ろし (tall,n=1.07) winston cone 22.7% 28%  20% 0.45  1.2

表 2: W20cm×H30cm
光学系集光率反射回数
主鏡ファネル平均値最高値最低値壁ファネル
平面鏡  winston (din=30) 26.1% 34%  18% 0.74  1.6
平面鏡 (tall)  winston (din=30) 25.2% 33%  17% 0.80  1.8
振り上げ振り下ろし  winston  30.0% 35%  24% 0.68  1.46
振り上げ振り下ろし (tall) winston (din=30) 30.7% 35% 28% 0.74  1.7
振り上げ振り下ろし (tall) winston (din=25) 31.9% 36%  27% 0.75  1.6
振り上げ振り下ろし (tall) winston (din=20) 30.1% 36%  26% 0.75  1.4

表 3: W20cm×H40cm
光学系集光率反射回数
主鏡ファネル平均値最高値最低値壁ファネル
平面鏡  winston (din=30) 20.2% 29%  12% 0.84  1.8
平面鏡 (tall)  winston (din=30) 19.3% 29%  12% 0.88  1.8
振り上げ振り下ろし  winston (din=30) 26.8% 31%  23% 0.78  1.69
振り上げ振り下ろし (tall) winston (din=30) 27.4% 30% 24% 0.81  1.8
振り上げ振り下ろし (tall) winston (din=25) 27.1% 31%  25% 0.81  1.6
振り上げ振り下ろし (tall) winston (din=20) 23.2% 27%  20% 0.83  1.4